# AEIA-MN: Evaluating the Robustness of Multimodal LLM-Powered Mobile Agents Against Active Environmental Injection Attacks

### v1

@@ Abstract
As researchers continuously optimize AI agents to perform tasks more effectively within operating systems, they often neglect to address the critical need for enabling these agents to identify "impostors" within the system. Through an analysis of the agents’ operating environment, we identified a potential threat: attackers can disguise their attack methods as environmental elements, injecting active disturbances into the agents’ execution process, thereby disrupting their decision-making. We define this type of attack as Active Environment Injection Attack (AEIA). Based on this, we propose AEIA-MN, an active environment injection attack scheme that exploits interaction vulnerabilities in the mobile operating system to evaluate the robustness of MLLM-based agents against such threats. Experimental results show that even advanced MLLMs are highly vulnerable to this attack, achieving a maximum attack success rate of 93% in the AndroidWorld benchmark.
AEIA-MN: Evaluating the Robustness of Multimodal LLM-Powered Mobile Agents Against Active Environmental Injection Attacks
Yurun Chen1,  Xueyu Hu1,  Keting Yin1,  Juncheng Li1,  Shengyu Zhang1 1Zhejiang University yurunchen123@gmail.com {huxueyu,yinkt,junchengli,sy_zhang}@zju.edu.cn
1 Introduction
@@
Recently, a growing body of research Durante et al. (2024); Nguyen et al. (2024); Hu et al. (2024a); Xi et al. (2025) on multimodal large language models (MLLMs) and commercial products Anthropic (2024); Apple Inc. (2024); Google (2024); OpenAI (2025) focused on (M)LLM-based agents has emerged, which we refer to as Operation System Agents (OS Agents). These agents leverage (M)LLMs for decision-making and utilize the built-in functionalities of operating systems (e.g., Windows, macOS, Android, iOS) to execute user-defined tasks. However, due to the inherent limitations of (M)LLMs, OS Agents are exposed to significant security threats.
Figure 1: The evaluation results of agents under attack from Active Environment Injection Attack (AEIA).
@@
The operating system environment contains various interference factors that could pose security risks to OS Agents (Wu et al., 2024a; Zhang et al., 2024; Wu et al., 2024b), yet research on active attack risks remains limited. Existing research (Ma et al., 2024; Liao et al., 2024) has introduced the concept of environment injection attacks, which typically affect the decision-making process of (M)LLMs by embedding adversarial content or inserting malicious code in webpages, with a focus on passive-trigger interference factors within the browser environment. However, there has been limited attention to the active attack risks faced by operating systems, creating a significant gap in OS Agent security. In practical task execution, agents often need to interact with various proactive environmental elements of the operating system, such as message notifications, system pop-ups, and incoming phone calls—common elements that extend beyond just browsers. Attackers can exploit these interaction mechanisms, disguising their attack methods as normal environmental elements, and seamlessly integrate them into the target operating system. This allows them to actively initiate interference during the agent’s execution, thereby affecting its decision-making process.
@@
Based on the threats outlined above, this paper further expands the concept of environmental injection attacks by introducing the notion of Active Environmental Injection Attack, named AEIA. We define this attack as: A malicious behavior that disguises attack vectors as environmental components through analysis of target operating system characteristics, and actively disrupts agent decision-making processes via specific system interaction mechanisms. The innovation of this attack paradigm lies in transcending traditional passive attack limitations, creating a novel attack surface tightly coupled with agent decision-making through dynamic environmental manipulation. We summarize the key characteristics of AEIA as follows: (i) Active injection: The attack should enable real-time modification of OS-level environmental parameters during agent operation. (ii) Process sensitivity: The effectiveness of the attack is highly reliant on the precise synchronization of the agent’s execution process. (iii) Environment integration: The attack techniques need to be integrated with the characteristics of a specific operating system environment for effective interaction.
@@
To demonstrate the effectiveness of AEIA, we propose AEIA-MN, an attack scheme that carries out Active Environmental Injection Attacks via Mobile Notifications. This approach leverages the unique characteristics of mobile environments by disguising the attack as a message notification, thus creating an attack strategy that exploits the interaction between operating system elements. Due to portability requirements, message notifications take up a larger portion of the screen on mobile devices compared to desktop devices, leading to greater interference with the decision-making process of MLLMs. Additionally, due to the pop-up feature of notifications, the attack can actively modify the device environment, in line with the characteristics of AEIA. Building on this, we design three distinct attack strategies targeting both the perception and reasoning phases based on the execution flow of the OS Agent. These strategies enable a comprehensive evaluation of the robustness of mobile OS agents based on different MLLMs against AEIA. Extensive experiments on two benchmarks validate the rationale and effectiveness of this attack, with attack success rates reaching up to 93%(AndroidWorld) and 84% (AppAgent), respectively. Overall, our contributions can be summarized as follows:
@@
  • We are the first to introduce the concept of Active Environment Injection Attack (AEIA), which offers a fresh perspective on the security research of OS Agents.
@@
  • Based on mobile agents, we design an active environmental injection attack scheme via mobile notifications named AEIA-MN that effectively evaluates the robustness of existing MLLM-based Agents against such attacks.
@@
  • We implemented a prototype of AEIA-MN and conducted extensive experiments to evaluate the security of various MLLM-based mobile agents when facing such attacks. The results indicate that current MLLMs have limited defensive capabilities against this attack.
2 Related Work
@@
The OS Agent refers to an agent tool that utilizes the capabilities of (M)LLMs to perform user-defined tasks within an operating system, based on native system operations. Notable web-based agents include SeeAct (Zheng et al., 2024), SeeClick (Cheng et al., 2024), and WebAgent (Gur et al., 2024), while mobile agents consist of InfiGUIAgent (Liu et al., 2025), AppAgent (Zhang et al., 2023), Mobile-Agent (Wang et al., 2024b), and Mobile-Agent-v2 (Wang et al., 2024a). Additionally, there are various other agents (Yan et al., 2023; Li et al., 2023; Wu et al., 2024c; Tan et al., 2024; Lee et al., 2024; Hoscilowicz et al., 2024; Deng et al., 2024; Hu et al., 2024b). These OS agents are often exposed to various security risks.
@@
Many studies on the security of OS agents have been conducted. Wu et al. (2024b) discovered a novel network threat called Web Indirect Prompt Injection (WIPI), which involves embedding natural language instructions in webpages to indirectly control web agents driven by large language models to execute malicious commands. Wu et al. (2024a) conducted a study shows that image perturbations can affect MLLMs to produce adversarial captioners, leading agents to pursue goals contrary to user intentions. Ma et al. (2024) highlighted the vulnerability of GUI agents to environmental disturbances and proposed an "Environment Injection" attack. Liao et al. (2024) introduced an environment injection attack scheme that injects malicious code in webpage to steal users’ personal identification information (PII). However, their work was limited to passive-triggered environmental injection attacks. Zhang et al. (2024) explored how to carry out pop-up window attacks on visual and language model (VLM)-based agents. Nevertheless, their discussion was confined to pop-up attacks in browser environments, missing broader research on environmental injection attacks. Yang et al. (2024) identified eight potential attack paths that agents on mobile devices might face but did not address active environmental injection attacks like those involving message notifications.
3 Methods
@@
In this section, we define the workflow of the OS Agent (Section 3.1), outlining the key stages and their interactions within the system. Furthermore, we provide a comprehensive explanation of the attack scheme AEIA-MN (Section 3.2).
3.1 Definition of the OS Agent Workflow
@@
To clarify the workflow of the OS Agent, the workflow parameters are formally defined as follows: $Action_{i}$ represents the action, and $State_{i}$ denotes the state during execution (with $State_{sys}$ representing the system state), where $i\in[0,N]$, with $i$ indicating the step required for the task and $N$ being the maximum number of steps needed to complete the task. $Action_{fin}$ represents the action indicating that the reasoning process is considered finished by the (M)LLMs. $State_{goal}$ denotes the goal state, $State_{mem}$ refers to the past state stored in memory, and $State_{fin}$ represents the final state after reasoning. $Reasoning(.)$ denotes the reasoning function, and $System(.)$ represents the system state function. We present the overall workflow of the OS Agent in Algorithm 1.
@@
3       Perception Stage: The OS Agent collects data from the OS. $State_{i}\leftarrow Perception(OS)$
@@
11The OS Agent finishes the task and obtains the final system state. $State_{fin}\leftarrow State_{i}$
Algorithm 1 Workflow of the OS Agent
Figure 2: The overview of our proposed attack scheme AEIA-MN.
3.2 Attack Scheme
We introduce AEIA-MN, an Active Environmental Injection Attack scheme that leverages mobile notifications to attack mobile agents on mobile devices. The objective of AEIA-MN is to induce a deviation in the agent’s decision-making from its intended goal, thereby enabling the evaluation of the model’s robustness against this attack. AEIA-MN encompasses three types of attacks: Adversarial Attack (Section 3.2.1), Reasoning Gap Attack (Section 3.2.2), and Combinatorial Attack (Section 3.2.3). We provide detailed information on the attack methods in AEIA-MN in Figure 2. Additionally, a comprehensive case study is presented in Appendix A, and the design of the mobile notifications is discussed in Appendix B.
3.2.1 Adversarial Attack
@@
Message Misleading. Attackers can craft notifications containing adversarial content, directly intervening in the Agent’s decision-making process through message-push channels. Such attacks exploit the vulnerabilities of natural language processing by embedding misleading instructions into the notification text using malicious means, thereby enticing the Agent to perform unintended actions.
UI Element Occlusion. The unique interface characteristics of mobile platforms create favorable conditions for occlusion attacks: (i) Spatial Proportion: Based on mainstream mobile and laptop devices, the average screen area covered by the notification bar reaches 13.4% (based on measurements from Android 13), significantly higher than the 1.2% on laptop platforms (based on measurements from a 14-inch MacBook display). (ii) Element Coverage: Notifications always float above the application interface, directly covering key interactive areas, such as the search bar (occurrence rate 92%) and functional buttons (an average of 3.2 per screen). This overlap can hinder the user’s ability to interact with these essential features.
It is noteworthy that the text display limitations of mobile devices necessitate a more concise approach to adversarial text information. According to the Android developer documentation Android Developers (2025), Android notification contant can display an average of approximately 40 characters (configured with the default font size on Android) when collapsed. As a result, this physical limitation forces attackers to construct adversarial texts with high information density.
@@
3.2.2 Reasoning Gap Attack
In addition to the Adversarial Attack, we identify a potential attack— the Reasoning Gap Attack. This attack exploits the systemic flaws of MLLM-based Agents to achieve unexpected state transitions. The decision-making process of (M)LLMs typically requires a certain reasoning time, which is acceptable in normal interactive dialogues. However, as an engineered tool, existing MLLM-based OS Agents continue to operate in a conversational interaction format, failing to adequately consider changes in device states during the reasoning period. Specifically, MLLMs experience significant time delays (approximately 0.5 to 5 seconds) from the initiation to the completion of reasoning, during which the system is in a state like Stop-The-World [^1] Stop-The-World typically refers to the process of pausing or stopping all activities at a specific moment, a common state during Java garbage collection and system maintenance., unable to respond to environmental changes or receive external inputs, creating a dangerous Reasoning Gap. This freezing of system state can lead to severe consequences in dynamic scenarios.
@@ 3.2.2 Reasoning Gap Attack
3.2.3 Combinatorial Attack
Given that attacks based on message notifications may vary depending on the execution phases of the OS Agent (e.g., deploying an Adversarial Attack during the perception phase and a Reasoning Gap Attack during the reasoning gap phase), attackers have the potential to combine these attacks. In AEIA-MN, attackers can utilize either an Adversarial Attack or a Reasoning Gap Attack individually, or combine both to form a more sophisticated and potent Combinatorial Attack. The core idea of the Combinatorial Attack is to amplify the disruptive effects by concurrently exploiting adversarial perturbations and reasoning gaps.
4 Evaluation
In this section, we implement a prototype of our attack scheme AEIA-MN and evaluate the robustness of different agents against the attack through extensive experiments. We first describe our experimental setup and metrics in Section 4.1 and 4.2, and then present and discuss the experimental results in Section 4.3, 4.4, and 4.5.

### v2

# Evaluating the Robustness of Multimodal Agents Against Active Environmental Injection Attacks

@@ Abstract.
As researchers continue to optimize AI agents for more effective task execution within operating systems, they often overlook a critical security concern: the ability of these agents to detect ”impostors” within their environment. Through an analysis of the agents’ operational context, we identify a significant threat—attackers can disguise malicious attacks as environmental elements, injecting active disturbances into the agents’ execution processes to manipulate their decision-making. We define this novel threat as the Active Environment Injection Attack (AEIA). Focusing on the interaction mechanisms of the Android OS, we conduct a risk assessment of AEIA and identify two critical security vulnerabilities: (1) Adversarial content injection in multimodal interaction interfaces, where attackers embed adversarial instructions within environmental elements to mislead agent decision-making; and (2) Reasoning gap vulnerabilities in the agent’s task execution process, which increase susceptibility to AEIA attacks during reasoning. To evaluate the impact of these vulnerabilities, we propose AEIA-MN, an attack scheme that exploits interaction vulnerabilities in mobile operating systems to assess the robustness of MLLM-based agents. Experimental results show that even advanced MLLMs are highly vulnerable to this attack, achieving a maximum attack success rate of 93% on the AndroidWorld benchmark by combining two vulnerabilities.
Multimodal Large Language Models, Agent Security, Operation System
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: the 33nd ACM International Conference on Multimedia (MM’25),; June 03–05, 2025; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Security and privacy Domain-specific security and privacy architectures
1. Introduction
@@
Recently, a growing body of research on multimodal large language models (MLLMs) (Durante et al., 2024; Nguyen et al., 2024; Hu et al., 2024a; Xi et al., 2025) and an increasing number of commercial products (Anthropic, 2024; Apple Inc., 2024; Google, 2024; OpenAI, 2025) focused on (M)LLM-based agents have emerged. These agents leverage (M)LLMs for decision-making and utilize built-in Operating System (OS) functionalities to execute user-defined tasks. However, due to the inherent limitations of (M)LLMs, these agents are vulnerable to significant security threats.
Figure 1. We evaluated the robustness of the agents under active attacks on environmental elements (notifications).
@@
Our research focuses on the threats posed by actively interactive environmental elements within operating systems (Figure 1 is an example). Previous studies (Ma et al., 2024; Liao et al., 2024) have introduced the concept of environment injection attacks, which typically affect the decision-making process of (M)LLMs by embedding adversarial content or inserting malicious code in webpages, with a focus on passive-trigger interference factors within the browser environment. However, there has been limited attention to the active attack risks faced by operating systems. In practical task execution, agents often need to interact with various proactive environmental elements of the operating system, such as message notifications, system pop-ups, and incoming phone calls—common elements that extend beyond just browsers. Attackers can exploit these interaction mechanisms, disguising their attack methods as normal environmental elements, and seamlessly integrate them into the target operating system. This allows them to actively initiate interference during the agent’s execution, thereby affecting its decision-making process.
@@
Based on the threats outlined above, we further expand the concept of environmental injection attacks by introducing the concept of Active Environmental Injection Attack, named AEIA. We define this attack as: a malicious behavior that disguises attack vectors as environmental elements by analyzing target operating system characteristics, then actively disrupts agent decision-making processes through specific system interaction mechanisms. The advantages of this attack concept lie in two aspects: first, it overcomes the limitation of passive triggering in environment injection attacks; second, it enables the active manipulation of environmental elements to design novel attack strategies that are tightly coupled with the execution process of agents. We summarize the key characteristics of AEIA as follows: (i) Active injection: The attack should enable real-time modification of OS-level environmental parameters during agent operation. (ii) Process sensitivity: The effectiveness of the attack is highly dependent on the precise synchronization of the agent’s execution process. (iii) Environment integration: The attack techniques need to be integrated with the characteristics of a specific operating system environment for effective interaction.
The goal of our work is to evaluate the robustness of agents against AEIA in Android OS. Existing research (Ma et al., 2024; Liao et al., 2024; Zhang et al., 2024) on environmental injection attacks has been limited to desktop environment and passive triggering methods, without exploring further threats related to actively interacting environmental elements. In this work, our key innovation lies in the further introduction of the novel attack concept AEIA and the extension of existing research to mobile OS. We explored the interaction system of the Android OS and discovered novel AEIA approaches based on mobile notifications. These approaches leverage the unique characteristics of mobile environments by disguising the attack as a message notification, thus creating an attack strategy that exploits the interaction between operating system elements.
@@
To evaluate security threat of the aforementioned AEIA attack in mobile environments, we propose AEIA-MN, an attack scheme that carries out Active Environmental Injection Attacks via Mobile Notifications. This attack scheme is designed around mobile notifications and the execution characteristics of agents, targeting both the perception and reasoning stages with three distinct attack strategies. These strategies enable a comprehensive evaluation of the robustness of agents based on different MLLMs against AEIA. Extensive experiments on two test benchmarks demonstrate that existing agents exhibit insufficient robustness against such attacks, with attack success rates reaching up to 93% (AndroidWorld) and 84% (AppAgent), respectively. Overall, our contributions in this work can be summarized as follows:
@@
  • To the best of our knowledge, we are the first to introduce the concept of Active Environment Injection Attack (AEIA), which provides a novel perspective on the security research of agents in operating systems.
@@
  • Based on the AEIA concept, we developed AEIA-MN, an active environment injection attack scheme leveraging mobile notifications, to effectively assess the robustness of existing MLLM-based agents against such attacks.
@@
  • We implemented a prototype of AEIA-MN and conducted extensive experiments to evaluate the security of various MLLM-based mobile agents when facing such attacks. The results indicate that current MLLM-based agents have limited defensive capabilities against this attack.
2. Related Work
@@
There has been extensive research on agents in both web and mobile environments. Notable web-based agents include SeeAct (Zheng et al., 2024), SeeClick (Cheng et al., 2024), and WebAgent (Gur et al., 2024), while mobile agents consist of InfiGUIAgent (Liu et al., 2025), AppAgent (Zhang et al., 2023), Mobile-Agent (Wang et al., 2024c), and Mobile-Agent-v2 (Wang et al., 2024b). Additionally, there are various other agents (Yan et al., 2023; Li et al., 2023; Wu et al., 2024a; Tan et al., 2024; Lee et al., 2024; Hoscilowicz et al., 2024; Deng et al., 2024; Hu et al., 2024b). Due to the complex design of operating systems, these agents are often exposed to various security risks.
@@
Many studies on the security of OS agents have been conducted. Wu et al. (2024c) discovered a novel network threat called Web Indirect Prompt Injection (WIPI), which involves embedding natural language instructions in webpages to indirectly control web agents driven by large language models to execute malicious commands. Wu et al. (2024b) conducted a study shows that image perturbations can affect MLLMs to produce adversarial captioners, leading agents to pursue goals contrary to user intentions. Ma et al. (2024) highlighted the vulnerability of GUI agents to environmental disturbances and proposed an ”Environment Injection” attack. Liao et al. (2024) introduced an environment injection attack scheme that injects malicious code in webpage to steal users’ personal identification information (PII). However, their work was limited to passive-triggered environmental injection attacks. Zhang et al. (2024) explored how to carry out pop-up window attacks on visual and language model (VLM)-based agents. Nevertheless, their discussion was confined to pop-up attacks in browser environments, missing broader research on environmental injection attacks. Yang et al. (2024) identified eight potential attack paths that agents on mobile devices might face but did not address active environmental injection attacks like those involving message notifications.
3. Methodology
@@
In this section, we define the workflow of the Agent (Section 3.1), outlining the key stages and their interactions within the system. Furthermore, we provide a comprehensive explanation of the attack scheme AEIA-MN (Section 3.2).
Figure 2. The overview of our proposed attack scheme AEIA-MN. We demonstrate the attack by adding the contact Hugo Pereira in the Contacts App (a task from the AndroidWorld benchmark). The left side illustrates the Agent’s workflow, which consists of Perception, Reasoning, and Action. The right side presents the details of the Adversarial Attack and Reasoning Gap Attack.
3.1. Definition of the Agent Workflow
@@
To clarify the workflow of the Agent, the workflow parameters are formally defined as follows: $Action_{i}$ represents the action, and $State_{i}$ denotes the state during execution, where $i\in[1,N]$, with $i$ indicating the $i$-th step required for the task and $N$ being the maximum number of steps needed to complete the task. In addition, $State_{sys}$ represents the system state. $Action_{fin}$ represents the action indicating that the reasoning process is considered finished by the MLLMs. $State_{goal}$ denotes the goal state, $State_{mem}$ refers to the past state stored in memory, and $State_{fin}$ represents the final state after reasoning. $Reasoning(.)$ denotes the reasoning function, and $System(.)$ represents the system state function. We present the overall workflow of the Agent in Algorithm 1.
@@
3       Perception Stage: The Agent collects data from the OS. $State_{i}\leftarrow Perception(OS)$
@@
11The Agent finishes the task and obtains the final system state. $State_{fin}\leftarrow State_{i}$
Algorithm 1 Workflow of the Agent
3.2. Attack Scheme
We introduce AEIA-MN, an Active Environmental Injection Attack scheme that leverages mobile notifications to attack mobile agents on mobile devices. The attack objective of AEIA-MN is to induce a deviation in the agent’s decision-making from its intended goal, thereby enabling the evaluation of the model’s robustness against this attack. We first presented the design of message notifications (Section 3.2.1). Then, we show two types of attacks and their combination encompassed by AEIA-MN: Adversarial Attack (Section 3.2.2), Reasoning Gap Attack (Section 3.2.3), and Combinatorial Attack (Section 3.2.4). We provide detailed information on the attack methods in AEIA-MN in Figure 2. Additionally, a comprehensive case study about AEIA-MN is discussed in supplementary material.
3.2.1. Notification Design
Message notifications on mobile devices are implemented in two main ways: Application(App)-based notifications and Short Message Service (SMS) notifications. App-based notifications lack a unified design standard and typically require users to download the app in advance. In the current era of advanced mobile security, malicious code in apps is easily detected by devices, limiting the applicability of this approach in many scenarios. In contrast, SMS-based notifications offer better compatibility on Android devices and usually require no special configuration. This makes SMS a potential vector for malicious attacks. Given these considerations, we chose to design message notifications based on SMS. In all the elements of SMS, we will use the text content of message notifications to implement adversarial attacks. In addition, the accessibility (a11y) tree of message notifications also reflects the adversarial content.
3.2.2. Adversarial Attack
@@
Message Misleading. Attackers can craft notifications containing adversarial content, directly intervening in the Agent’s decision-making process through message-push channels. Such attacks exploit the vulnerabilities of MLLMs by embedding misleading instructions into the notification text using malicious means, thereby enticing the Agent to perform unintended actions.
UI Element Occlusion. The unique interface characteristics of mobile platforms create favorable conditions for occlusion attacks: (i) Spatial Proportion: Based on mainstream mobile and laptop devices, the average screen area covered by the notification bar reaches 13.4% (based on measurements from Android 13), significantly higher than the 1.2% on laptop platforms (based on measurements from a 14-inch MacBook display). (ii) Element Coverage: Notifications always float above the application interface, directly covering key interactive areas, such as the search bar (occurrence rate 92%) and functional buttons (an average of 3.2 per screen). This overlap can hinder the agent’s ability to interact with these essential features in the mobile OS.
It is noteworthy that the text display limitations of mobile devices necessitate a more concise approach to adversarial text information. According to the Android developer documentation (Android Developers, 2025), Android notification contant can display an average of approximately 40 characters (configured with the default font size on Android) when collapsed. As a result, this physical limitation forces attackers to construct adversarial texts with high information density.
@@
3.2.3. Reasoning Gap Attack
In addition to the Adversarial Attack, we identify a potential attack targeting agents in the OS—the Reasoning Gap Attack. This attack exploits the systemic flaws of MLLM-based Agents to achieve unexpected state transitions. The decision-making process of (M)LLMs typically requires a certain reasoning time, which is acceptable in normal interactive dialogues. However, as an engineered tool, existing MLLM-based Agents still operate in a conversational interaction format in the operating system, failing to adequately consider changes in device states during the reasoning period. Notably, MLLMs exhibit significant time delays—ranging from approximately 3 to 12 seconds (with detailed measurement methods provided in the supplementary materials)—between the initiation and completion of reasoning. During this interval, the system enters a state like Stop-The-World [^1] Stop-The-World is a term commonly used to describe a system state where all application threads are halted simultaneously, often observed during Java garbage collection or certain system maintenance operations., unable to respond to environmental changes or receive external inputs, creating a dangerous Reasoning Gap.
This freezing of system state (Reasoning Gap) can lead to severe consequences in dynamic scenarios: when critical changes occur in environmental parameters during the decision gap, the Agent will make decisions based on outdated state information, resulting in actions that do not align with the current environment. Adversaries can change the system state immediately by utilizing message notifications within a specific time window, causing the Agent to enter an unexpected state after executing operations based on outdated decisions. This attack has two threat characteristics: (i) Time-sensitive: it maximizes the success rate of attacks through precise timing control; (ii) Generalizability: it is applicable to all MLLM system architectures based on synchronous interaction paradigms.
@@ 3.2.3. Reasoning Gap Attack
3.2.4. Combinatorial Attack
Given that attacks based on message notifications may vary depending on the execution phases of the OS Agent (e.g., deploying an Adversarial Attack during the perception phase and a Reasoning Gap Attack during the reasoning gap phase), attackers have the potential to combine these attacks. In AEIA-MN, attackers can utilize either an Adversarial Attack or a Reasoning Gap Attack individually, or combine both to form the Combinatorial Attack. The core idea of the Combinatorial Attack is to amplify the disruptive effects by concurrently exploiting adversarial perturbations and reasoning gaps.
Table 1. The difference between the agents in input data.